Ticket Number: HEA-NOC/20080516-2 Ticket Status: CLOSE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20080516 11:16 UTC+1 Problem Start: 20080516 11:30 UTC+1
Ticket Closed: 20080516 12:44 UTC+1 Problem End: 20080516 12:30 UTC+1
After a successful migration to HEAnet's new IP network, the old links to
the old IP network need to be removed.
There will be no outage to services. Traffic is routing to new IP network.
20080516 11:14 GL Commencing work
20080516 12:44 GL Completed. Closing ticket
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20080516-2
HEAnet Limited. Registered in Ireland, No. 275301.